# Balanced Similarity with Auxiliary Prompts: Towards Alleviating Text-to-Image Retrieval Bias for CLIP in Zero-shot Learning

### v1

@@ Abstract
CLIP has the ability to align texts and images and is nearly the most frequently used foundation model in cross-modal zero-shot learning. However, our experimental findings reveal that CLIP suffers from a bias in text-to-image retrieval, resulting in a decrease in CLIP’s zero-shot learning performance. We analytically discover that the bias partly arises from the imbalanced range of similarity scores obtained by CLIP. Accordingly, we propose a Balanced Similarity with Auxiliary Prompts (BSAP) to mitigate the text-to-image retrieval bias of CLIP. Specifically, our BSAP designs auxiliary prompts for CLIP to calculate multiple similarity scores for the retrieval images and then normalizes the scores between each image and the given query text as well as our auxiliary prompts to obtain balanced similarity scores. The balanced similarity score of the given query text is used for the final retrieval. In addition, we attempt to adopt a hybrid similarity that combines our BSAP with the original similarity of CLIP to obtain a more robust outcome. Extensive experiments on two typical zero-shot learning tasks, i.e., Referring Expression Comprehension (REC) and Referring Image Segmentation (RIS), are conducted to demonstrate the effectiveness of our BSAP. Specifically, when using the val dataset of RefCOCO in REC, BSAP increases CLIP’s performance by 20.6%.
1 Introduction
Figure 1: (a) The illustration of the text-to-image retrieval bias of CLIP. The text-to-image retrieval bias is caused by the ranges of the similarity scores. (b) The performance comparison with and without BSAP of CLIP-based method [39] when matching image regions with given texts under three datasets.
@@ 1 Introduction
Training models for specific tasks necessitates extensive data collection, a process that can be remarkably time-consuming and expensive. In contrast, zero-shot learning [33, 3] eliminates the need for the training step by transferring existing models and making slight modifications to the architecture. Considering its efficiency, zero-shot learning has attracted increasing attention recently.
@@
Recent advancements in foundation models [5, 15, 25, 26, 13] have revolutionized the fields of computer vision and natural language processing and have brought feasible solutions for zero-shot learning considering their capability and generalizability. Among all foundation models, CLIP [34] is the most representative one that can align texts and images and is nearly the most commonly used one that improves cross-modal zero-shot tasks, including object detection [9], semantic segmentation [52], image captioning [40], and image-text retrieval [2].
However, we discovered that CLIP suffers from the text-to-image retrieval bias, which significantly decreases the zero-shot learning performance. Here, the text-to-image retrieval bias refers to the fact that CLIP would obtain suboptimal results when asking CLIP to figure out which image perfectly matches one given query text among several candidate images. Naturally, the bias is commonly believed to be caused by the low distinguishing ability of CLIP to falsely identify the contents of images, and it is nearly impossible to solve such a problem in zero-shot learning without finetuning to improve the CLIP.
In contrast, in this paper, we analytically found that in some situations, the CLIP actually could understand the semantic texts and images, and some text-to-image retrieval bias is experimentally proven to be caused based on the imbalanced range of similarity scores calculated by CLIP. As shown in Figure 1 (a), CLIP can classify the categories of dog and person images; however, CLIP misidentifies the dog image as the person text because the similarity score between the dog image and person text is higher than the similarity score between person image and person text. Accordingly, we argue that such type of text-to-image retrieval bias of CLIP can be mitigated if we can balance the similarity scores in zero-shot learning.
@@
In light of the above analysis, we propose a balanced similarity with auxiliary prompts (BSAP) for CLIP to conduct text-to-image retrieval, which mitigates the text-to-image retrieval bias caused by the imbalanced range of similarity scores. Specifically, our BSAP first generates auxiliary prompts for the CLIP to conduct the text-to-image retrieval, and then normalizes the scores between each image and the given query text as well as our auxiliary prompts to obtain a balanced similarity score, which is used in the final retrieval. As shown in Figure 1 (b), with BSAP, the text-to-image retrieval bias of CLIP-based method [39] is alleviated. Note that our auxiliary prompts are pre-defined, and the BSAP requires no training process. In addition, we further design a hybrid similarity that combines our BSAP with the original similarity of CLIP to obtain a more robust outcome. We conducted extensive experiments on two typical zero-shot learning tasks, i.e., Referring Expression Comprehension (REC) and Referring Image Segmentation (RIS). Three datasets, including RefCOCOg [17, 29], RefCOCO+ [30], and RefCOCO [30], and several state-of-the-art CLIP-based methods [39, 49] are adopted. The experimental results demonstrate that BSAP improves the CLIP-based methods. Specifically, the results in the REC task of all datasets were increased by an average point of $5.92\%$.
@@
  • We discover the text-to-image retrieval bias in CLIP, which significantly decreases CLIP’s performance in zero-shot learning. Moreover, we analytically prove that some biases are caused by the imbalanced range of similarity scores calculated by the CLIP.
  • We propose a balanced similarity with auxiliary prompts (BSAP) for CLIP to mitigate the text-to-image retrieval bias. By introducing the auxiliary prompts, we normalize the original similarity scores to alleviate the imbalanced range problems. The experimental results on two typical zero-shot learning tasks, i.e., REC and RIS, demonstrate the capability of our BSAP.
2 Related Work
Zero-shot Learning.
@@ 2 Related Work
Foundation Models.
@@
Foundation Models [32] are models trained on large corpora of data that, at a very large scale, can generalize to new tasks without any task-specific finetuning. The primary method considered in this study is CLIP [34]. Our focus on CLIP is rooted in its pre-training on a dataset comprising 400 million image-text pairs gathered from the web, enabling impressive zero-shot image classification performance across various visual domains. CLIP incorporates a dedicated image encoder, either based on the ResNet architecture [12] or a visual transformer [8], alongside a distinct text transformer. The primary utilization involved the RN50x16 and ViTB/32 versions of CLIP. The limited discriminative capabilities of CLIP are also observed [37, 45]. This paper focuses on the text-to-image retrieval bias of CLIP.
Referring Expression Comprehension and Referring Expression Segmentation.
@@
Referring expression comprehension (REC) [48] and referring expression segmentation (RES) [28] require text-to-image retrieval and are typical zero-shot visual grounding [53] tasks. In recent advancements in zero-shot learning for visual grounding, the prevalent approach often employs the transferable CLIP model [39, 49]. The central module involves transforming the task into a text-to-image retrieval task, complemented by additional auxiliary design modules. Six methods utilizing CLIP for zero-shot REC and RIS tasks are introduced. For REC task, Colorful Prompt Tuning (CPT) [47] applies different colors to various proposals and uses a masked language prompt. The pre-trained masked language model [41] like CLIP calculates the color with the highest probability. GradCAM [36] is a commonly used gradient visualization technique, employed to understand which parts of the input image are most crucial for the model’s predictions in computer vision tasks. ReCLIP [39] consists of two components. The first component is a region-scoring method that isolates object proposals through cropping and blurring before passing them to CLIP. The second component is a spatial relation resolver, capable of handling various spatial relationships proposed by the authors. Our approach involved conducting experiments on the REC and RES tasks.
@@
For RIS task, Region Token [23] is a technique employed in adapting CLIP. Similar to the adaptation process for CLIP, it utilizes region tokens for each mask suggestion in all Transformer layers of the CLIP visual encoder, instead of using superpixels. Finally, it calculates the cosine similarity between the class label of each mask proposal and the text features of CLIP, selecting the mask with the highest score. Cropping [22, 11, 44] is a commonly used method that leverages CLIP to extract mask or bounding box features in a range of zero-shot dense prediction tasks. Zero-shot-RIS [49] introduces a mask-oriented visual encoder capturing both global and local contextual information from the input image. By utilizing instance masks obtained from off-the-shelf mask proposal techniques, this method can perform fine-grained instance-level groundings segmentation. Additionally, the author introduces a global-local text encoder, where global features capture complex sentence-level semantics of the entire input expression, while local features focus on target noun phrases extracted by a dependency parser.
3 Text-to-Image Retrieval Bias for CLIP
(a) Examples of Category Misclassifications by Reclip.
Figure 2: Examples and analyses of bias in CLIP on text-to-image retrieval tasks. The blue bounding boxes represent the Ground Truth (GT), while the red bounding boxes depict the predicted results from ReClip.
(b) The Imbalanced Range of the Similarity Scores Obtained by CLIP.
(a) Examples of Category Misclassifications by Reclip.
Figure 2: Examples and analyses of bias in CLIP on text-to-image retrieval tasks. The blue bounding boxes represent the Ground Truth (GT), while the red bounding boxes depict the predicted results from ReClip.
@@ 3 Text-to-Image Retrieval Bias for CLIP
A text-to-image retrieval bias exists in CLIP. We similarly observed the same issue in the CLIP-base model and proved it use REC tasks in three datasets, i.e., RefCOCOg [17, 29], RefCOCO+ [30], and RefCOCO [30] validation, as illustrations. We found that using the Reclip model based on CLIP, nearly $15\%$ of the text cannot be aligned with the correct corresponding images in all the mentioned datasets. We present several examples in Figure 1(a) 1(a). One intuitive reason is that the CLIP cannot identify the semantic contents of these images, leading to the wrong region image selection for the query text. Without training or fine-tuning, it is almost impossible to improve CLIP’s distinguishing ability in zero-shot learning.
(a) Results for caption “person”.
(b) Results for caption “dog”.
Figure 3: The illustration of the similarity scores generated by the CLIP model, in which 100 images of dog and 100 images of human are selected. In the graphs, blue dots represent the result for dog images and red dots represent the result for human images. (a) Similarity processed by CLIP for the text caption “person” with 100 images of dog and 100 images of human. The results for dog images are sorted from large to small and the results for human images are sorted from small to large. (b) Similarity processed by CLIP for the caption text “dog” with 100 images of dog and 100 images of human. The results for the dog images are sorted from small to large and the results for human images are sorted from large to small.
@@
However, it was discovered another problem that was generally ignored would also cause the text-to-image retrieval bias, i.e., the imbalanced range of the similarity scores obtained by the CLIP. One case is shown in Figure 1(a) 1(b). It was observed that the smallest score of Region-1 is still higher than the highest score of Region-2. Naturally, Region-1 would always be selected. To further analyze the imbalanced range problem, we randomly selected 100 images of dogs and 100 images of persons. To our surprise, when asking whether the images belong to text descriptions of a dog or a person, CLIP can obtain a 99.5% accuracy. This means that the CLIP has a strong ability to understand the contents of dog and person images and text descriptions. Nevertheless, the Figure 3 shows the similarity scores between all images and the text descriptions of the dog and person obtained using CLIP. The similarity scores for 100 dog images paired with the caption “person” are arranged in descending order, and the similarity results for 100 human images paired with the caption “person” are arranged in ascending order. We observe that the distribution of dog similarity scores is mixed with the distribution of person similarity scores when using person descriptions with images. A significant portion of dog images are higher than those of person images; in other words, dog images will be selected as the match with the caption “person” in most situations. The imbalanced similarity score ranges harm the identifying process.
4 The Proposed Method
@@
In this section, we design a balanced similarity with auxiliary prompts (BSAP) to alleviate the text-to-image retrieval bias for CLIP in zero-shot learning. Firstly, we present the details to obtain the balanced similarity score and then give an explanation of how to generate auxiliary prompts under zero-shot learning. In addition, we introduce a hybrid similarity, which combines our BSAP and original similarity.
4.1 Balanced Similarity with Auxiliary Prompts
@@
Figure 4: The process of BSAP to get balanced similarity with auxiliary prompts.
@@
In light of the analysis in Section 3, the diverse ranges of similarity scores would lead to the text-to-image retrieval bias. Intuitively, if we can balance all similarity scores in the same space, the problem could be mitigated. Inspired by the normalization strategies [35, 1], we propose a balanced similarity with auxiliary prompts (BSAP).
@@
As shown in Figure 4, there are several auxiliary text prompts that could be used as calibration. Similarity scores can be adjusted with these auxiliary prompts. Specifically, for m-th image with one textual referring expression and auxiliary prompts, the similarity scores of the image are calculated as follows:
@@
where $\text{Sim}_{*}$ refers to the similarity score of CLIP, $\mathbf{V}_{m}$ refers to the visual feature of the m-th image, $\mathbf{T}_{\text{ref}}$ refers to the textual embeddings of query text, and $\mathbf{T}_{\text{ap}_{i}}$ refers to the textual embeddings of auxiliary prompts, $i\in[A]$ and $A$ is the number of auxiliary prompts.
Then, we normalize the scores and obtain the balanced similarity score. The process is defined as follows:
@@
where $\mathbf{BS}$ is the final balanced similarity score. The process is similar to a SoftMax function. Note that other normalization strategies can also be adopted, such as feature scaling, we select SoftMax because it experimentally obtains sufficient improvements. If the Auxiliary Prompts are sufficient, we can obtain reliable balanced scores, since the normalization would balance the range of similarity scores of different objects in a similar space. The final prediction is determined as follows:
@@
where $\text{BS}_{m}$ refers to the balanced similarity score of $m$-th image.
4.2 Auxiliary Prompt Generation
One key component of our BSAP is how to obtain the Auxiliary Prompts under zero-shot learning settings since there are no reference templates. We design our Auxiliary Prompts from two aspects: contents and templates.

### v2

# Towards Alleviating Text-to-Image Retrieval Hallucination for CLIP in Zero-shot Learning

@@ Abstract
Pretrained cross-modal models, for instance, the most representative CLIP, have recently led to a boom in using pre-trained models for cross-modal zero-shot tasks, considering the generalization properties. However, we analytically discover that CLIP suffers from the text-to-image retrieval hallucination, adversely limiting its capabilities under zero-shot learning: CLIP would select the image with the highest score when asked to figure out which image perfectly matches one given query text among several candidate images even though CLIP knows contents in the image. Accordingly, we propose a Balanced Score with Auxiliary Prompts (BSAP) [^1] Related codes will be released upon acceptance. to mitigate the CLIP’s text-to-image retrieval hallucination under zero-shot learning. Specifically, we first design auxiliary prompts to provide multiple reference outcomes for every single image retrieval, then the outcomes derived from each retrieved image in conjunction with the target text are normalized to obtain the final similarity, which alleviates hallucinations in the model. Additionally, we can merge CLIP’s original results and BSAP to obtain a more robust hybrid outcome (BSAP-H). Extensive experiments on two typical zero-shot learning tasks, i.e., Referring Expression Comprehension (REC) and Referring Image Segmentation (RIS), are conducted to demonstrate the effectiveness of our BSAP. Specifically, when evaluated on the validation dataset of RefCOCO in REC, BSAP increases CLIP’s performance by 20.6%. Further, we validate that our strategy could be applied in other types of pretrained cross-modal models, such as ALBEF and BLIP.
Index Terms:
Pretrained cross-modal models, CLIP, text-to-image retrieval hallucination, zero-shot learning
I Introduction
@@
Training models for specific tasks necessitate extensive data collection, a process that can be remarkably time-consuming and expensive. In contrast, zero-shot learning [1, 2] eliminates the need for the training step by transferring existing models and making slight modifications to the architecture. Considering its efficiency, zero-shot learning has attracted increasing attention recently.
@@
Recent advancements in foundation models or pretrained models [3, 4] have revolutionized the fields of computer vision and natural language processing and have brought feasible solutions for zero-shot learning considering their capability and generalizability. Among all foundation models, CLIP [5] is the most representative one that can align texts and images and is nearly the most commonly used one that improves cross-modal zero-shot tasks, including object detection [6], semantic segmentation [7], image captioning [8], and image-text retrieval [9].
However, we discovered that CLIP suffers from a type of Text-to-Image Retrieval hallucination, which significantly decreases its zero-shot learning performance. Here, the Text-to-Image Retrieval hallucination refers to the fact that CLIP would obtain a wrong result when asking CLIP to figure out which image perfectly matches one given query text among several candidate images even though CLIP understands the contents in the image. We analytically found that in some situations, the CLIP actually can understand the semantic texts and images, and it has been demonstrated that a portion of the hallucinations in CLIP is due to the substantial variance in the range distribution of results computed by CLIP for the same text groups across different images. For instance, as shown in Fig. LABEL:fig:first (a), CLIP can identify categories of dog and person in the image. However, CLIP misclassified the dog image as the person text because the similarity score between the dog image and the person text is higher than the similarity score between the person image and the person text. Therefore, we posit that if we can address the issue of imbalanced range distribution in outcomes of multi-modal large foundation models during zero-shot learning, we may be able to mitigate some of the hallucinations in models like CLIP.
@@
In light of the aforementioned analysis, we propose to use Balanced Scores with Auxiliary Prompts (BSAP) for CLIP in text-to-image retrieval, thereby mitigating hallucinations stemming from the imbalanced distribution of the outcome range. Specifically, taking the CLIP model as an example, our BSAP initially generates a series of auxiliary result sets for CLIP. It then standardizes the scores between each image and given query texts against our auxiliary score sets to achieve a balanced distribution of outcome ranges for the final retrieval. As shown in Fig. LABEL:fig:first (b), with BSAP, the hallucination caused by the imbalanced distribution in CLIP-based method [10] is alleviated. Note that our auxiliary prompts are pre-defined, and the BSAP requires no training process. In addition, we further obtain hybrid results using BSAP (BSAP-H) that combine balanced results using BSAP with the original results created by Vision-Language Models (VLMs) to obtain more robust final results. We conducted extensive experiments on two typical zero-shot learning tasks, i.e., Referring Expression Comprehension (REC) and Referring Image Segmentation (RIS). Three datasets, including RefCOCOg [11, 12], RefCOCO+ [13], and RefCOCO [13], and several state-of-the-art CLIP-based methods [10, 14] are adopted. The experimental results demonstrate that BSAP improves the CLIP-based methods and some other VLMs. Specifically, the results in the REC task of all datasets were increased by an average point of $5.92\%$.
@@
  • We analytically discover that pretrained cross-modal models, for example CLIP, suffer from text-to-image retrieval hallucination, which significantly degrades CLIP’s performance in zero-shot learning.
  • We propose Balanced Scores with Auxiliary Prompts (BSAP) to mitigate hallucinations caused by the imbalanced distribution of results. By calculating a set of reference outcomes, we normalize the original results in conjunction with the reference set to balance the raw results.
  • Experimental results on two typical zero-shot learning tasks, namely REC and RIS, demonstrate the capability of our BSAP and BSAP-H. Both tasks have achieved state-of-the-art performance. Additionally, we have briefly tested our BSAP on other vision-language models and the image-to-text retrieval task, and it has also brought about a noticeable improvement.
II Related Works
II-A Zero-shot Learning
@@
II-B Foundation Models
Foundation Models [19] are models trained on large corpora of data that, at a very large scale, can generalize to new tasks without any task-specific finetuning.
@@ II-B Foundation Models
The primary method considered in this study is CLIP [5]. Our focus on CLIP is rooted in its pre-training on a dataset comprising 400 million image-text pairs gathered from the web, enabling impressive zero-shot image classification performance across various visual domains. CLIP incorporates a dedicated image encoder, either based on the ResNet architecture [20] or a visual transformer [21], alongside a distinct text transformer. The primary utilization involved the RN50x16 and ViTB/32 versions of CLIP. The limited discriminative capabilities of CLIP are also observed [22]. The paper primarily explores that CLIP the same as MLLMs, also exhibits hallucinations characteristic of such models.
II-C Hallucination in Vision-Language Models
Hallucinations [23, 24] denote the emergence of text that is irrelevant, factually inaccurate, or nonsensical within a specific context—a significant challenge faced by contemporary large foundational models. This issue may stem from overfitting to particular patterns observed in the training data, a deficiency in grasping real-world factuality, or a failure to contextualize the input provided appropriately [25]. Our research has uncovered that hallucinations are not confined to text-generation tasks. They also manifest in the text-to-image retrieval task domain, which is our study’s focus. In this task, discrepancies arise when searching for images that correspond to textual descriptions using VLMs. Nonetheless, the models demonstrate an aptitude for accurately classifying both erroneous and accurate images. This proficiency in classification suggests that VLMs possess the capability to discern between correct and incorrect images. Consequently, our findings have shed light on the phenomenon of hallucinations within the context of large models engaged in text-to-image retrieval tasks.
II-D Referring Expression Comprehension and Referring Expression Segmentation
@@
Referring expression comprehension (REC) [26] and referring expression segmentation (RES) [27] require text-to-image retrieval and are typical zero-shot visual grounding [28] tasks. In recent advancements in zero-shot learning for visual grounding, the prevalent approach often employs the transferable CLIP model [10, 14]. The central module involves transforming the task into a text-to-image retrieval task, complemented by additional auxiliary design modules. Six methods utilizing CLIP for zero-shot REC and RIS tasks are introduced. For the REC task, Colorful Prompt Tuning (CPT) [29] applies different colors to various proposals and uses a masked language prompt. The pre-trained masked language model [30] like CLIP calculates the color with the highest probability. GradCAM [31] is a commonly used gradient visualization technique, employed to understand which parts of the input image are most crucial for the model’s predictions in computer vision tasks. ReCLIP [10] consists of two components. The first component is a region-scoring method that isolates object proposals through cropping and blurring before passing them to CLIP. The second component is a spatial relation resolver, capable of handling various spatial relationships proposed by the authors. Our approach involved conducting experiments on REC and RES tasks.
@@
For the RIS task, Region Token [32] is a technique employed in adapting CLIP. Similar to the adaptation process for CLIP, it utilizes region tokens for each mask suggestion in all Transformer layers of the CLIP visual encoder, instead of using superpixels. Finally, it calculates the cosine similarity between the class label of each mask proposal and the text features of CLIP, selecting the mask with the highest score. Cropping [33] is a commonly used method that leverages CLIP to extract mask or bounding box features in a range of zero-shot dense prediction tasks. Zero-shot-RIS [14] introduces a mask-oriented visual encoder capturing both global and local contextual information from the input image. This method can perform fine-grained instance-level groundings segmentation by utilizing instance masks obtained from off-the-shelf mask proposal techniques. Additionally, the author introduces a global-local text encoder, where global features capture complex sentence-level semantics of the entire input expression, while local features focus on target noun phrases extracted by a dependency parser.
III Hallucination in CLIP
@@
Hallucination is discovered in CLIP during the text-to-image retrieval task. We observed the issue in the CLIP-base model and proved it use REC tasks in three datasets, i.e., RefCOCOg [11, 12], RefCOCO+ [13], and RefCOCO [13] validation, as illustrations. We found that using the Reclip model based on CLIP, nearly $15\%$ of the text cannot be aligned with the correct corresponding images in all the mentioned datasets. We present several examples in Fig. LABEL:fig:imbalanced (a). One intuitive reason is that the CLIP misclassifies the semantic contents of these images, leading to the wrong region image selection for the query text. Without training or fine-tuning, it is almost impossible to improve CLIP’s distinguishing ability in zero-shot learning.
@@
It was discovered the imbalanced range of the results obtained by the CLIP, which was generally ignored, was one of the reasons that led to hallucination when performing text-to-image retrieval. One case is shown in Fig. LABEL:fig:imbalanced (b). It was observed that the smallest score of Region-1 is still higher than the highest score of Region-2. Naturally, Region-1 would always be selected. To further analyze the imbalanced range problem, we randomly selected 100 images of dogs and 100 images of persons. To our surprise, when asking whether the images belong to text descriptions of a dog or a person, CLIP can obtain a 99.5% accuracy. This means that the CLIP has a strong ability to understand the contents of dog and person images and text descriptions. Nevertheless, the Fig. LABEL:fig:cosine_similarity shows outcomes between all images and the text descriptions of the dog and person obtained using CLIP. The outcomes for 100 dog images paired with the caption “person” are arranged in descending order, and the outcomes for 100 human images paired with the caption “person” are arranged in ascending order. We observe that the distribution of dog results is mixed with the distribution of person results when using person descriptions with images. A significant portion of dog images are higher than those of person images; in other words, dog images will be selected as the match with the caption “person” in most situations. The imbalanced result ranges lead to hallucinations.
IV The Proposed Method
@@
In this section, we design Balanced Scores with Auxiliary Prompts (BSAP) to alleviate the hallucinations for CLIP in zero-shot learning. Firstly, we present the details to obtain balanced results and then give an explanation of how to generate auxiliary prompts under zero-shot learning. In addition, we introduce hybrid results using BSAP (BSAP-H), which combines our BSAP and original outcomes.
IV-A Balanced Scores with Auxiliary Prompts
@@
Figure 2: The process of BSAP to get balanced results with auxiliary prompts.
@@
In light of the analysis in Section III, the diverse ranges of results would lead to hallucinations. Intuitively, if we can balance all outcomes in the same space, the problem could be mitigated. Inspired by the normalization strategies [34, 35], we propose Balanced Scores with Auxiliary Prompts (BSAP).
@@
As shown in Fig. 2, there are several auxiliary text prompts that could be used as calibration. Similarity scores can be adjusted with these auxiliary prompts. Specifically, for $m$-th image with one textual referring expression and auxiliary prompts, similarity scores of the image are calculated as follows:
@@
where $\text{Sim}_{*}$ refers to similarity scores of CLIP, $\mathbf{V}_{m}$ refers to the visual feature of the $m$-th image, $\mathbf{T}_{\text{ref}}$ refers to textual embeddings of query text, and $\mathbf{T}_{\text{ap}_{i}}$ refers to textual embeddings of auxiliary prompts, $i\in[A]$ and $A$ is the number of auxiliary prompts.
Then, we normalize scores and obtain balanced similarity scores. The process is defined as follows:
@@
where $\mathbf{BS}$ is final balanced similarity scores. The process is similar to a SoftMax function. Note that other normalization strategies can also be adopted, such as feature scaling. We select SoftMax because it experimentally obtains sufficient improvements. If the auxiliary prompts are sufficient, we can obtain reliable balanced scores, since the normalization would balance the range of similarity scores of different objects in a similar space. The final prediction is determined as follows:
@@
where $\text{BS}_{m}$ refers to the balanced result score of $m$-th image.
IV-B Auxiliary Prompts Generation
One key component of our balanced outcome scores is how to obtain the auxiliary prompts under zero-shot learning settings since there are no reference templates. We design our auxiliary prompts from two aspects: contents and templates.